The time required for an elevator to lift a weight varies jointly as the weight as the weight and distance through which it is to be lifted and inversely as the power of the motor. It takes 30 seconds for a 10 horsepower motor to lift 100 pounds through 50 ft. What size of the motor is required to lift 800 pounds in 40 seconds through 40 feet?
1. What type variation is involved in this problem?
2. What previously learned concepts are needed in solving the problem?
3. What is the mathematical equation of the given problem?
4. What is the unknown in the problem?
5.How will you solve the given problem involving combined variation?

Answer by Boreal(15235)

t=k*w*d/p
30=k*100*50/10
k=0.06
use that in next step
40=0.06*800*40/p
40p=1920
p=48 hp